OpenStudy (anonymous):

Solve for x. –2(6 – x) = 3x A. –3 B. 12 C. 3 D. –12

OpenStudy (shamil98):

Distributive property. -2(6-x)=3x -12 + 2x = 3x -12 = x

OpenStudy (anonymous):

thank you @shamil98 ;*